In a random sample of six mobile​ devices, the mean repair cost was ​$80.00 and the standard deviation was ​$12.00. Assume the population is normally distributed and use a​ t-distribution to find the margin of error and construct a 95​% confidence interval for the population mean. Interpret the results. The 95​% confidence interval for the population mean mu is
Answer #1

Solution: The provided Yo formation is sis follasuks Saraple Bean -80 Sample size sample standard devisition s= 12 confidence

= (80-12:593, 80+ 120593) CI = (67-403, 92.593) Interpretation Therefore based on the data provided the 95% confidence intern